# Superfluid Band Theory for the Rod Phase in the Magnetized Inner Crust Matter: Entrainment, Spin-orbit, Spin-triplet Pairing

### v1

@@ 4.1 On Entrainment
Figure 1: Calculated neutron effective masses as a function of magnetic-field strength. The red solid line indicates the results when the magnetic direction is along $x$-axis, while the green dash-dotted line is dedicated to $z$-axis result.
@@ 4.1.1 Effective Mass
Table 1 summarizes the calculated conduction number density and effective mass for several magnetic field strengths. Calculations are performed for two magnetic-field orientations: along the $x$-axis, perpendicular to the nuclear rods, and along the $z$-axis, parallel to the rods. The dimensionless magnetic-field strengths $B_{\star}=100$ and $1000$ correspond to approximately $4.4\times 10^{15}$ G and $4.4\times 10^{16}$ G, respectively, both of which are within the expected range for magnetars. This table indicates that, in the absence of a magnetic field, the neutron effective mass is approximately $m^{\star}_{n}/m_{n}\simeq 1.5$, indicating a clear entrainment regime. This result is consistent with the band-theoretical calculations by Carter et al. Carter et al. (2005), whereas it contrasts with previous one-dimensional studies, including Ref. Kashiwaba and Nakatsukasa (2019); Sekizawa et al. (2022) and our earlier work Yoshimura and Sekizawa (2024), where anti-entrainment was observed. However, it remains unclear how these results might change when superfluidity is explicitly included or when the effective mass is derived via time-evolution calculations; thus, this point remains open for further discussion. Regarding the magnetic-field dependence, the entrainment effect remains almost unchanged for moderate field strengths around $B_{\star}=100$. In contrast, at $B_{\star}=1000$, the effective mass exceeds twice the bare neutron mass, corresponding to an enhancement by a factor of approximately 1.5 compared to the zero-field case. Figure 1 shows the effective mass as a function of magnetic-field strength, demonstrating a smooth increase from $m_{n}^{\star}/m_{n}\simeq 1.4$ to values exceeding 2.0 over the explored range. Notably, this behavior is essentially independent of the magnetic-field orientation, with similar trends observed for both $x$- and $z$-directions. As mentioned earlier, $B_{\star}=100$$1000$ corresponds to approximately $10^{15}$$10^{16}$ G in cgs units, a magnitude expected to exist within magnetars. These findings suggest that conventional frameworks for astrophysical dynamics, may not be applicable to magnetars in their current form. Consequently, this implies the necessity for more detailed future investigations into the internal structure and dynamics of these systems in the presence of strong magnetic fields.
Figure 2: The energy band diagram along the Brillouin zone path of the square lattice. The left figure shows the results of no magnetic-field case, whereas the right is for $B_{z}=1000B_{\star}$ case. The black dashed line in two figures indicate the chemical potentials of neutrons.
@@ 4.1.2 Band Diagram
Figure 2 presents the calculated neutron energy band structures along a representative path in the Brillouin zone for the square lattice, both in the absence of a magnetic field and in the presence of a magnetic field of $B_{z}=1000B_{\star}$. In this figure, a clear enhancement of entrainment can be observed under strong magnetic fields, which can be traced back to qualitative changes in the band structure. In the zero-field case, the energy bands are doubly degenerate owing to Kramers degeneracy. When a magnetic field is applied, this degeneracy is lifted by the Zeeman effect, leading to a proliferation of distinct energy bands near the Fermi surface. The resulting dense spectrum promotes band flattening through level repulsion, which in turn enhances the neutron effective mass. A simple estimate indicates that the Zeeman splitting at $B_{\star}=1000$ is of the order of $0.5$ MeV, exceeding the typical inter-band spacing near the Fermi level. This leads to a highly intricate band structure in that energy region and provides a microscopic explanation for the substantial increase in the effective mass. Thus, by examining the detailed band structure, we explicitly demonstrate how strong magnetic fields modify macroscopic transport properties such as the neutron effective mass.
4.2 On Spin-orbit
In this section, we present and discuss the results obtained from Hartree-Fock-Bogoliubov (HFB) calculations that explicitly incorporate pairing correlations. For the sake of computational simplicity, band-structure effects are neglected in this subsection by setting $N_{k_{x}}=N_{k_{y}}=1$. The influence of band-structure effects will be examined separately in a subsequent section.
@@ 4.2 On Spin-orbit
Figure 3 plots the averaged pairing gap of the spin-singlet neutron superfluid as a function of magnetic field strength. The results are displayed for magnetic fields applied along the $x$- and $z$-axes, and for cases with and without the spin–orbit term in the energy density functional. This figure indicates that the pairing gap remains a nearly constant up to a certain threshold of the magnetic field strength, beyond which, it decreases abruptly and vanishes, signaling a transition to the normal phase. This qualitative behavior is robust and does not depend on the magnetic-field direction or on the inclusion of the spin-orbit interaction.
@@
Figure 4 plots the magnitude of the neutron spin-polarization as a function of the magnetic field strength. Results for magnetic fields applied along the $x$- and $z$-axes are shown in separate panels, with calculations performed both with and without the spin–orbit term. Note that the vertical scales differ between the two panels, being of order $10^{-3}$ for the $x$-axis case and $10^{-2}$ for the $z$-axis case. Two key features can be drawn from these results. In the fitst place, in the absence of the spin-orbit interaction, the spin-polarization is entirely determined by the magnetic-field direction. No polarization appears for magnetic fields applied along the $x$-axis, whereas for fields along the $z$-axis the polarization remains zero up to a threshold field strength of approximately $5000,B_{\star}$, above which it emerges abruptly. This behavior is naturally understood from the definition of the spin-polarization in terms of the $z$-component of the spin density [Eq. (64)] and from the fact that the Zeeman coupling acts directly on this component. This result is also consistent with our previous findings Yoshimura and Sekizawa (2025). In the second place, the inclusion of the spin-orbit interaction qualitatively modifies the polarization behaviour. Namely, a finite, albeit small, spin-polarization emerges even when the magnetic field is applied along the $x$-axis, with an onset at much lower field strengths of the order of $1000\,B_{\star}$. This behavior was not observed in previous one-dimensional calculations, where the spin-orbit interaction yields no contribution owing to its rotational nature. The present results therefore demonstrate, for the first time, that the interplay between dimensionality and spin-orbit coupling can induce spin-polarization in superfluid neutron matter under magnetic fields.
4.3 On Spin-triplet pairing
@@ 4.3.1 Spin-triplet EDF
In this section, we discuss the computational results concerning spin-triplet superfluid condensation. Figures 5 and 6 show the magnitude of the rank components of the neutron spin-triplet pairing as functions of the magnetic-field strengths. Figure 5 corresponds to calculations without the spin-triplet pairing EDF, while Fig. 6 presents the results obtained when this interaction is included. By comparing these two figures, three conclusions can be drawn. In the first place, in both calculations, rank-0 component emerges abruptly once the magnetic-field strength exceeds a critical value, reaching magnitudes of the order of $10^{-3}$. This threshold takes place around $B_{\star}=5000$, which coinsides with the critical magnetic field for the appearance of spin polarization. This correspondence indicates that the emergence of the rank-0 component is closely linked to the spin-polarized phase, consistent with previous studies highlighting the relationship between polarized neutron matter and the $^{3}P_{0}$ pairing order parameter Tajima et al. (2023). In the second place, the rank-2 component never appears in the absense of the spin-triplet pairing EDF, whereas it acquires finite values whenever the corresponding interaction channel is included, largely independent of the magnetic-field strength. This result demonstrates that the $^{3}P_{2}$ pairing component is driven primarily by the presense of the spin-triplet interaction itself, in contrast to the $^{3}P_{0}$ component, which is induced by spin polarization. In the third place, the rank-1 component remains negligibly small in all cases, irrespective of the magnetic-field direction, strength, or the inclusion of the spin-triplet pairing EDF. Since the rank-1 component corresponds to the vector part of the pairing spin current, which is forbidden by spatial inversion symmetry, its suppression can be attributed to the inversion symmetry of the two-dimensional rod phase around the $z$-axis.
@@ 4.3.2 Band Structure Effects
Thus far, the results have been obtained without explicitly including band wavenumbers. We now examine how band-structure effects influence spin-triplet pairing.
Table 2 summarizes the calculated rank-0 and rank-2 components with and without band-structure effects for representative magnetic-field strengths. From this table, it is found that the qualitative dependence of each component on the magnetic-field remains essentially unchanged when band structure effects are included, although their magnitudes are quantitatively modified. An exception is found for the $^{3}P_{2}$ component under magnetic fields applied along the $x$-axis, where band-structure effects lead to an enhancement by approximately one order of magnitude. Nevertheless, this enhancement does not alter the overall conclusion: the emergence of spin-triplet superfluidity is governed primarily by spin-polarization and by the presence of specific interaction channels, and remains qualitatively robust against crystalline band-structure effects.
@@ 5 SUMMARY and PROSPECTS
In this work, we have extended the framework of the superfluid band theory, developed in our previous studies, to the two-dimensional rod phase of the neutron star inner crust. For describing the spin-triplet neutron superfluidity, the $S=1$ component of the anomalous density is incorporated in the superfluid DFT, on the same footing with the effects of finite magnetic-fields. Within this unified framework, we have systematically investigated spin-polarization and pairing phases by selectively switching on and off the spin-orbit interaction and spin-triplet pairing EDF.
Our main findings can be summarized as follows. In the first place, on the magnetic-fields with $10^{16}$ G order, the effective masses of neutrons increase by a factor of around $1.5$ compared to no magnetized cases. This enhancement originates from a drastic modification of the band structure when the Zeeman splitting becomes comparable to the inter-level spacing. The result indicates that the conventional application of band theory to crustal matter may become questionable in magnetars, and calls for a more careful microscopic treatment of neutron dynamics under such extreme conditions. In the second place, we have examined the response of spin polarization and pairing gaps to magnetic fields, with particular attention to the role of the spin-orbit interaction. While the pairing gap has been found to depend primarily on the magnitude of the magnetic field, the spin-orbit interaction enables the emergence of weakly spin-polarized phases even for moderate field strengths and non-aligned field directions. Such effects have not been captured in previous superfluid band calculations, which were restricted to one-dimensional geometries where spin-orbit contributions vanish. Our spin-dependent HFB calculations in the two-dimensional rod phase thus reveal, for the first time, the nontrivial interplay between dimensionality, spin-orbit coupling, and magnetic fields in crustal matter. In the third place, we have systematically investigated the emergence of spin-triplet superfluid components. The results have demonstrated that the rank-0 component arises solely from magnetic-field-induced spin polarization, whereas the rank-2 component is generated by specific spin-triplet pairing interactions. By comparing calculations with and without band-structure effects, we further showed that these superfluid properties remain qualitatively robust in crystalline environments, although their magnitudes are quantitatively modified. Self-consistent nuclear calculations of neutron spin-triplet pairing remain scarce, and their application to neutron-star crust matter is pioneered in the present study. These findings provide a microscopic basis for future investigations of astronomical simulations involving unconventional superfluid phases.
@@ ACKNOWLEDGEMENT
We would like to appreciate Kenichi Yoshida (RCNP, Osaka University) and Nobuo Hinohara (Tsukuba University) for the kind introduction of the formalism of the spin-triplet pairing. The author K.Y. acknowledges Shunsuke Yasunaga (Institute of Science Tokyo) who checked the formulation of this work from a theoretical point of view. This work is supported by JSPS Research Fellow, Grant No. JP24KJ1110, as well as JSPS Grantin-Aid for Scientific Research, Grants No. JP23K03410 and No. JP23K25864. This work mainly used the computational resources of TSUBAME4.0 at Institute of Science Tokyo, through the HPCI System Project, Project ID: hp230180, hp240183, and hp250097. Additionally this work used computational resources of the Yukawa-21 supercomputer at Yukawa Institute for Theoretical Physics (YITP), Kyoto University. This work partly used computational resource of Miyabi-C provided by Multidisciplinary Cooperative Research Program (MCRP) in Center for Computational Science, University of Tsukuba, Project ID: xg25i043. This work partly used computational resource of Fugaku provided by RIKEN through the HPCI System Research Project (Project ID: hp250288).

### v2

@@ 4.1 Entrainment effects
Figure 1: Calculated neutron effective masses as a function of magnetic-field strength. The red solid (green dash-dotted) line corresponds to the results when the magnetic direction is along $x$-axis ($z$-axis) perpendicular (parallel) to the rods.
@@ 4.1.1 Effective mass
Table LABEL:tab:Entrainment summarizes the calculated conduction number density and effective mass for several magnetic field strengths. Calculations are performed for two magnetic-field orientations: along the $x$-axis, perpendicular to the nuclear rods, and along the $z$-axis, parallel to the rods. The dimensionless magnetic-field strengths $B_{\star}=100$ and $1000$ correspond to approximately $4.4\times 10^{15}$ G and $4.4\times 10^{16}$ G, respectively, both of which are within the expected range for magnetars. This table indicates that, in the absence of magnetic field, the neutron effective mass is approximately $m^{\star}_{n}/m_{n}\simeq 1.5$, indicating a clear entrainment regime. This result is consistent with the band-theoretical calculations by Carter et al. Carter et al. (2005), whereas it contrasts with previous one-dimensional studies, including Ref. Kashiwaba and Nakatsukasa (2019); Sekizawa et al. (2022) and our earlier work Yoshimura and Sekizawa (2024), where anti-entrainment was observed. However, it remains unclear how these results might change when superfluidity is explicitly included or when the effective mass is derived via time-evolution calculations; thus, this point remains open for further discussion. Regarding the magnetic-field dependence, the entrainment effect remains almost unchanged for moderate field strengths around $B_{\star}=100$. In contrast, at $B_{\star}=1000$, the effective mass exceeds twice the bare neutron mass, corresponding to an enhancement by a factor of approximately 1.5 compared to the zero magnetic-field case. Figure 1 shows the effective mass as a function of magnetic-field strength, demonstrating a smooth increase from $m_{n}^{\star}/m_{n}\simeq 1.4$ to values exceeding 2.0 over the explored range. Notably, this behavior is essentially independent of the magnetic-field orientation, with similar trends observed for both $x$- and $z$-directions. As mentioned earlier, $B_{\star}=100$$1000$ corresponds to approximately $10^{15}$$10^{16}$ G in cgs units, a magnitude expected to exist within magnetars. These findings suggest that conventional frameworks for astrophysical dynamics, may not be applicable to magnetars in their current form. Consequently, this implies the necessity for more detailed future investigations into the internal structure and dynamics of these systems in the presence of strong magnetic fields.
Figure 2: The energy band diagram along a representative path in the Brillouin zone of the square lattice. The left figure shows the results of zero magnetic-field case, whereas the right is for $B_{z}=1000B_{\star}$ case. The black dashed line in two figures indicate the chemical potentials of neutrons.
@@ 4.1.2 Band diagram
Figure 2 presents the calculated band structures of neutrons along a representative path in the Brillouin zone for the square lattice, both in the absence of a magnetic field (left panel) and in the presence of a magnetic field of $B_{z}=1000B_{\star}$ (right panel). In this figure, a clear enhancement of entrainment can be observed under strong magnetic fields, which can be traced back to qualitative changes in the band structure. In the zero magnetic-field case, the energy bands are doubly degenerate owing to Kramers degeneracy. When a magnetic field is applied, this degeneracy is lifted by the Zeeman effect, leading to proliferation of distinct energy bands near the Fermi surface. The resulting dense spectrum promotes band flattening through level repulsion, which in turn enhances the neutron effective mass. A simple estimate indicates that the Zeeman splitting at $B_{\star}=1000$ is of the order of $0.5$ MeV, exceeding the typical band spacing near the Fermi level. This leads to a highly intricate band structure in that energy region and provides a microscopic explanation for the substantial increase in the effective mass. Thus, by examining the detailed band structure, we explicitly demonstrate how strong magnetic fields modify macroscopic transport properties such as the neutron effective mass.
4.2 Spin-orbit couplings
In this section, we present and discuss the results obtained from HFB calculations that explicitly incorporate pairing correlations. For the sake of computational simplicity, band-structure effects are neglected in this subsection by setting $N_{k_{x}}=N_{k_{y}}=1$. The influence of band-structure effects will be examined separately in a subsequent section.
@@ 4.2 Spin-orbit couplings
Figure 3 plots the averaged pairing gap of the spin-singlet neutron superfluid as a function of magnetic field strength. The results are displayed for magnetic fields applied along the $x$- and $z$-axes, and for cases with and without the spin-orbit term in the energy density functional. This figure indicates that the pairing gap remains nearly constant up to a certain threshold of the magnetic field strength, beyond which, it decreases abruptly and vanishes, signaling a transition to the normal phase. This qualitative behavior is robust and does not depend on the magnetic-field direction or on the inclusion of the spin-orbit interaction.
@@
Figure 4 plots the magnitude of the neutron spin-polarization as a function of the magnetic field strength. Results for magnetic fields applied along the $x$- and $z$-axes are shown in separate panels, with calculations performed both with and without the spin-orbit term. Note that the vertical scales differ between the two panels, being of order $10^{-3}$ for the $x$-axis case and $10^{-2}$ for the $z$-axis case. Two key features can be drawn from these results. In the fitst place, in the absence of the spin-orbit interaction, the spin-polarization is entirely determined by the magnetic-field direction. No polarization appears for magnetic fields applied along the $x$-axis, whereas for fields along the $z$-axis the polarization remains zero up to a threshold field strength of approximately $5000\,B_{\star}$, above which it emerges abruptly. This behavior is naturally understood from the definition of the spin-polarization in terms of the $z$-component of the spin density [Eq. (64)] and from the fact that the Zeeman coupling acts directly on this component. This result is also consistent with our previous findings Yoshimura and Sekizawa (2025). In the second place, the inclusion of the spin-orbit interaction qualitatively modifies the polarization behaviour. Namely, a finite, albeit small, spin-polarization emerges even when the magnetic field is applied along the $x$-axis, with an onset at much lower field strengths of the order of $1000\,B_{\star}$. This behavior was not observed in previous one-dimensional calculations, where the spin-orbit interaction yields no contribution in the slab phase. The present results therefore demonstrate, for the first time, that the interplay between dimensionality and spin-orbit coupling can induce spin-polarization in superfluid neutron matter under magnetic fields.
4.3 Spin-triplet pairing
@@ 4.3.1 Spin-triplet EDF
In this section, we discuss the computational results concerning spin-triplet superfluid condensation. Figures 5 and 6 show the magnitude of the rank components of the neutron spin-triplet pairing as functions of the magnetic-field strengths. Figure 5 corresponds to calculations without the spin-triplet pairing EDF, while Fig. 6 presents the results obtained when this interaction is included. By comparing these two figures, three conclusions can be drawn. In the first place, in both calculations, the rank-0 component emerges abruptly once the magnetic-field strength exceeds a critical value, reaching magnitudes of the order of $10^{-3}$. This threshold takes place around $B_{\star}=5000$, which coinsides with the critical magnetic field for the appearance of spin polarization. This correspondence indicates that the emergence of the rank-0 component is closely linked to the spin-polarized phase, consistent with previous studies highlighting the relationship between polarized neutron matter and the $^{3}\text{P}_{0}$ pairing order parameter Tajima et al. (2023). In the second place, the rank-2 component never appears in the absense of the spin-triplet pairing EDF, whereas it acquires finite values whenever the corresponding interaction channel is included, largely independent of the magnetic-field strength. This result demonstrates that the $^{3}\text{P}_{2}$ pairing component is driven primarily by the presense of the spin-triplet interaction itself, in contrast to the $^{3}\text{P}_{0}$ component, which is induced by spin polarization. In the third place, the rank-1 component remains negligibly small in all cases, irrespective of the magnetic-field direction, strength, or the inclusion of the spin-triplet pairing EDF. Since the rank-1 component corresponds to the vector part of the pairing spin current, which is forbidden by spatial inversion symmetry, its suppression can be attributed to the inversion symmetry of the two-dimensional rod phase around the $z$-axis.
@@ 4.3.2 Band structure effects
Thus far, the results have been obtained without explicitly including Bloch wavenumbers. We now examine how band-structure effects influence spin-triplet pairing.
Table LABEL:tab:3PJ_band summarizes the calculated rank-0 and rank-2 components with and without band-structure effects for representative magnetic-field strengths. From this table, it is found that the qualitative dependence of each component on the magnetic-field remains essentially unchanged when band structure effects are included, although their magnitudes are quantitatively modified. An exception is found for the $^{3}\text{P}_{2}$ component under magnetic fields applied along the $x$-axis, where band-structure effects lead to an enhancement by approximately one order of magnitude. Nevertheless, this enhancement does not alter the overall conclusion: the emergence of spin-triplet superfluidity is governed primarily by spin-polarization and by the presence of specific interaction channels, and remains qualitatively robust against crystalline band-structure effects.
@@ 5 SUMMARY AND PROSPECTS
In this work, we have extended the framework of the superfluid band theory, developed in our previous studies, to the two-dimensional rod phase of the neutron star inner crust. Moreover, we have implemented spin-triplet pairing into the self-consistent superfluid band theory. For describing the spin-triplet neutron superfluidity, the $S=1$ component of the anomalous density is incorporated in the superfluid DFT, on the same footing with the effects of finite magnetic-fields. Within this unified framework, we have systematically investigated spin-polarization and pairing phases by selectively switching on and off the spin-orbit interaction and spin-triplet pairing EDF.
Our main findings can be summarized as follows. In the first place, on the magnetic-fields on the order of $10^{16}$ G, the effective masses of neutrons increase by a factor of around $1.5$ compared to no magnetized cases. This enhancement originates from a sizable modification of the band structure when the Zeeman splitting becomes comparable to the inter-level spacing. The result indicates that the conventional application of band theory to crustal matter may become questionable in magnetars, and calls for a more careful microscopic treatment of neutron dynamics under such extreme conditions. In the second place, we have examined the response of spin polarization and pairing gaps to magnetic fields, with particular attention to the role of the spin-orbit interaction. While the pairing gap has been found to depend primarily on the magnitude of the magnetic field, the spin-orbit interaction enables the emergence of weakly spin-polarized phases even for moderate field strengths and non-aligned field directions. Such effects have not been captured in previous superfluid band calculations, which were restricted to one-dimensional geometries where spin-orbit contributions vanish. Our spin-dependent HFB calculations in the two-dimensional rod phase thus reveal, for the first time, the nontrivial interplay between dimensionality, spin-orbit coupling, and magnetic fields in crustal matter. In the third place, we have systematically investigated the emergence of spin-triplet superfluid components. The results have demonstrated that the rank-0 component arises solely from magnetic-field-induced spin polarization, whereas the rank-2 component is generated by specific spin-triplet pairing interactions. By comparing calculations with and without band-structure effects, we further showed that these superfluid properties remain qualitatively robust in crystalline environments, although their magnitudes are quantitatively modified. Self-consistent nuclear calculations of neutron spin-triplet pairing remain scarce, and their application to neutron-star crust matter is pioneered in the present study. These findings provide a microscopic basis for future investigations of astronomical simulations involving unconventional superfluid phases.
@@ ACKNOWLEDGEMENT
We would like to appreciate Kenichi Yoshida (RCNP, Osaka University) and Nobuo Hinohara (Tsukuba University) for the kind introduction of the formalism of the spin-triplet pairing. The author K.Y. acknowledges Shunsuke Yasunaga (Institute of Science Tokyo) who checked the formulation of this work from a theoretical point of view. This work is supported by JSPS Research Fellow, Grant No. JP24KJ1110, as well as JSPS Grantin-Aid for Scientific Research, Grants No. JP23K03410, No. JP23K25864, and No. JP25H01269 This work mainly used the computational resources of TSUBAME4.0 at Institute of Science Tokyo, through the HPCI System Project, Project ID: hp230180, hp240183, and hp250097. Additionally this work used computational resources of the Yukawa-21 supercomputer at Yukawa Institute for Theoretical Physics (YITP), Kyoto University. This work partly used computational resource of Miyabi-C provided by Multidisciplinary Cooperative Research Program (MCRP) in Center for Computational Science, University of Tsukuba, Project ID: xg25i043. This work partly used computational resource of Fugaku provided by RIKEN through the HPCI System Research Project (Project ID: hp250288).